In a right triangle, if the legs are x feet and 3x+3 feet, and the hypotenuse is 25 feet, then solve for x by using the quadratic formula.
Answer by fractalier(2101)

The Pythagorean Theorem is
a^2 + b^2 = c^2
If we substitute in we have
x^2 + (3x + 3)^2 = 25^2 and
x^2 + 9x^2 + 18x + 9 = 625
10x^2 + 18x - 616 = 0
5x^2 + 9x - 308 = 0
(5x + 44)(x - 7) = 0
x = -44/5 or x = 7
Since distance can't be negative,
x = 7